OpenStudy (anonymous):

someone help please y=5x-6 solve for x (show work)

OpenStudy (azureilai):

x=(y+6)/5

OpenStudy (azureilai):

I can't actually get a numerical value answer unless there is some other piece of information.

OpenStudy (anonymous):

y=5x-6 y+6=5x(-6+6) (y+6)/5=5x/5 x=(y+6)/5